OpenStudy (anonymous):

Find the nth term of the sequence C^2, C^5, C^8 been thinking for this for about an hour now.. :/

OpenStudy (johnweldon1993):

Well as you can see...each time the power goes up by 3...c^2 to c^5 is a change of c^3 We know that to add exponents we multiply them...so to get from c^2 to c^5 we do c^2 * c^3 =c^(3 + 2) = c^5 So...from each term..you multiply by c^3 For nth-term = n-1* c^3 Example...for the 4th term... 4 - 1 = 3rd term = c^8 * c^3 = c^11 So on and so forth